40m QIL Cryo_Lab CTN SUS_Lab CAML OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log, Page 188 of 357  Not logged in ELOG logo
ID Date Author Type Category Subject
  8518   Wed May 1 10:42:35 2013 SteveUpdateLSCcleanup

 

 Optics from the car were placed into glass door cabinet E0

  8517   Wed May 1 00:05:03 2013 KojiUpdateLSCMore stable lock of PRMI (REF33I and AS55Q)

[Jenne Koji]

- Today the spots were moving more than the usual. The OPLEV screens showed that the spots are too much off from the center.

- Each vertex OPLEVs were checked and OPLEV wonderland was discovered: Other than the usual misalignment of the spots,
it was found that PRM/ITMX/ITMY beams were clipped somewhere in the paths, BS/PRM oplevs had many loose components
including the input lenses (they are still clamped by a single dog clamp THIS SHOULD BE FIXED ASAP).

- On the ITMY table there were so many stray optics. They were removed and put on the wagon next to the ITMY table.
THIS SHOULD BE CLEARED ON THE WEDNESDAY CLEANING SESSION.

- During this OPLEV session, LSCoffset nulling was run.

- After the OPLEV session, the locking became really instantaneous. We wonder which of the OPLEV cleaning, LSC offset nulling,
and the usual seismic activity decay in the evening was effective to make it better.

- Initially the lock was attempted with REFL33I/Q and some ~10sec lock streches were obtained. During this lock,
  the optical gain of AS55Q was measured in relative to REFL33Q. In deed they were calibrated to be the same
  gain at the input matrix.

- After the MICH signal source was switched to AS55Q, the lock streches became more regular and the minutes long.
We precisely tuned the phase of AS55 and REFL55 in terms of the differential excitation of ITMX/Y using lockin (FREQ 250, AMP 1000).

- We noticed that the AS port spot with AS55Q MICH was darker than the REFL33Q MICH. This suggests the existence of residual offset
in REFL33Q. In deed we observed +30cnt offset in REFL33Q when the PRMI is locked with AS55Q MICH.

- Phases and relative gains of the signals were as follows:

PRCL: REFL33I 1.00 =REFL55I +0.4
MICH: AS55Q 29deg x1.00 = REFL33Q -14deg x1.00 = REFL55Q 118deg 0.03?

- We tried to lock PRMI with AS55Q. The acquisition was not as easy as that with REFL33I. This might be from the saturation of the
REFL55I signal. This configuration should tested with different whitening gain. Handing off using the input matrix went well once the
lock was obtained by REFL33I.

- Handing off from AS55Q to REFL55Q was not successful.

- At the end of the session, Jenne told me that the POP PD still has a large diameter beam. (and a steering mirror with a peculiar reflection angle.)
==> THIS SHOULD BE FIXED ASAP
because the normalization factor can be too much susceptible to the misalignment of the spot.

- The configuration of the filters:

PRCL FM3/4/5/6 G=+0.05 / NORM 0.04 POP110I
MICH FM4/5 G=-5.00 / NORM 0.01 POP110I (or none)

Screenshot.png

  8516   Tue Apr 30 23:17:25 2013 JenneUpdateLSCPRCL LSC filters copied to CARM bank temporarily

Koji is working on PRMI locking with different photodiodes, and rather than typing different numbers into the input matrix, it is more convenient to just be able to click on/off buttons for different filter banks.  So, the CARM filter bank in the LSC model is currently being borrowed as a secondary PRCL filter bank.  I have copied all of the current PRCL filters over to the CARM filter bank. 

Just for reference, although we have not yet used CARM for CARM, the previous filters were the "default" set, +6dB, 0:1, 1:5, 1:50, 1000:10, RG3.2, RG16.5, RG24, empty, empty.  These are currently the same in the DARM and MC filter banks, so we can copy them back over in the future.

  8515   Tue Apr 30 23:04:23 2013 JenneConfigurationRF SystemOnly 4 25m cables ordered

I have found in the depths of the elog the (original?) list of fibers and lengths that were decided upon:  elog 6535.

In Suresh's elog, we were assuming that POP22 & POP110 would be served by a single PD.  This is still the nominal plan, although we (Rana is maybe still thinking about this in the back of his head?) think that it might not be feasible.  Riju and I were hoping to put a 4th fiber in the tubing so that we wouldn't have to add it later if POP22 & POP110 are eventually 2 separate PDs.  Anyhow, for now, all we have available are 3 fibers for the POX table, so that is what was installed this afternoon.

  8514   Tue Apr 30 22:40:57 2013 JenneUpdateSUSoplev XY-plots reflect new calibration

Back when Gabriele was here, he and I implemented online calibration of the oplevs, into microradians.  A consequence of this is that all of the XY-plots on the medm screens were too small. 

I have gone through all the screens that I could think of (SUS_SINGLE, SUS_SINGLE_OPTLEV_SERVO, OPLEV_MASTER, OPLEV_SUMMARY, OPLEV_SUMMARY_SMALL_SCALE, IFO_OVERVIEW) and made the range of the XY-plots +/- 100, rather than the old scale of +/-1. 

I have also added red boxes behind the numbers on many (but not yet all) of these screens, so that you can see when (a) the oplev sum is too low, or (b) either the pit or yaw value is over 50 microradians. 

While I was putzing around on the IFO overview screen, I also made the oplev sum numbers clickable, with the related display being that optic's oplev servo screen.

  8513   Tue Apr 30 21:24:15 2013 JenneConfigurationRF SystemPOX fiber laying

Nice work.  That was a lot of effort, but having done it so nicely will definitely pay off, since it is now much harder to break the fibers.

2 small issues:  In your attachment 3, I see a coil of fiber just outside the POX table.  I thought Koji had asked that all spare coiled-up length of fiber always be at the splitter side.  Also, in securing the plastic tubing as it comes down near the PSL table, you have zip-tied the tubing to the PSL table.  Since that is a space that we need to access to align the Xgreen beatnote stuff, please disconnect that zip tie, and secure the tubing on the north side somewhere, underneath the AP table, rather than the PSL table (when you look closer, you may notice that no cables in that bundle are attached to the PSL side at the bottom, for this same reason). 

  8512   Tue Apr 30 19:39:14 2013 RijuConfigurationRF SystemPD frequency response

Today I have rerouted the fibers on POX table. The aim was to lay it overhead through the plastic pipe. A pipe ~50ft (~15.5m) long was taken for this purpose. I disconnected the two 25m long fibers for POP55 and POX11 PD (those had been already routed) from both of their ends - i.e. from the POX table and also from 1x16 splitter. Jenne and Koji suggested that we may have another two PDs ( POP22 and POP110) on POX table in future. So we used another 25m long fiber for these two (POP22/POP110). We could not use two fibers for these two since we have only four 25m long fibers and one of them we need for POY11 PD on POY table. Jenne and me put the three fibers inside the pipe using a copper tube. The tube then was put on the overhead rack, Manasa helped me to do it. The fiber ends were finally laid on the POX table at one end and connected to the 1x16 splitter at the other end.

The corresponding photos are attached herewith.

  8511   Tue Apr 30 12:25:23 2013 ChloeUpdate QPD

Annalisa and I met yesterday and fixed the voltage regulator on the breadboard so the QPD circuit is working. We will meet with Eric on Thursday to determine the course of action with measurements.

  8510   Tue Apr 30 10:54:35 2013 JenneUpdateSUSETMX restored

I'm not sure why or when it was tripped, but I have restored the ETMX watchdog.

  8509   Mon Apr 29 23:02:48 2013 KojiConfigurationLSCQuestons

Q. How much Schnupp asymmetry we want in order to improve the signal ratio between PRCL/MICH in REFL ports?

Q. How much can we increase Schnupp asymmetry in the practical constraints?

Q. How PRCL/MICH ratio is different the REFL ports?
=> My modeling (many years ago) shows the ratio of {115, 51, 26, 23} for REFL{11, 33, 55, 165}.
These numbers should be confirmed by modern simulation of the 40m with updated parameters.
I should definitely use 55MHz but also prepare better 165MHz too.

Q. How the TT/PRM motions are affecting the lock stability? How can we quantify this effect? How can we mitigate this issue?

Q. Can we somehow change the sensing matrix by shifting the modulation frequency?

Q. Is normalization by POP22 or POP110 actually working well?
=> Time series measurement of error signals & servo inputs

  8508   Mon Apr 29 22:13:41 2013 KojiUpdateLSCLocking with ASDC

Today the locking was not as easy as that was last Friday.
So I tried something new. Today Rana talked about the ASDC locking with POPDC normalization.
This technique was tried. (This is somewhat similar to DC readout.)

PRCL
Signal source: REFL33I / Normalization POP110I x 0.04 / Trigger POP110I 20up 3down, otherwise  untouched from Friday locking
Servo: input matrix 1.00 -> PRCL Servo FM3/4/5/6 Always ON G=+0.06
Actuator: output matrix 1.00 -> PRM

MICH
Signal source: ASDC Offset -109.5 (nominal of the day -49.5) / Normalization POPDC x 1.00 / Trigger POP110I 20up 3down
Servo: input matrix 1.00 -> MICH Servo FM5 Always On G=+10000
ActuaroL output matrix -1.00 -> ITMX / +1.00 -> ITMY

Observation

- POP110I was ~120 during the lock (cf 170 on Friday). So there is some small leakage from the dark port.

- Lock was easier when FM4 of the MICH loop was turned off.

- During the lock horizontal motion of the intracavity mode was visible as usual.

Screenshot.png

  8507   Mon Apr 29 18:53:03 2013 JenneUpdateElectronics1pps timing fiber to OMC rack may be bent

While helping Riju out this afternoon, I noticed that the timing fiber that goes to the OMC rack (near the AP table) was bent, and is now possibly kinked, after the installation of the fiber splitter box. 

The fiber was hanging from the back of the rack, and had been strain relieved.  However, the path that the fiber was taking is now occupied by the fiber splitter for the RF PD diagnostic stuff.  So, the installation of the fiber splitter box put the old timing fiber under tension, causing the fiber to be bent at a little over 90 degrees, since it was pulled tightly against the corner of the splitter's front panel. 

I adjusted the strain relief so that the fiber is loose again, although there is still a bit of a kink that you can feel.  Things (for now) seem to be working, since the 1pps light on the front of the box at the top of the OMC rack is still blinking happily, indicating that the 1pps is still getting there. 

We are not using most of the stuff in that rack right now, but if we have problems in the future, we should check out the fiber to make sure it is still good.

  8506   Mon Apr 29 17:26:22 2013 RijuConfigurationRF SystemPD frequency response

 Today I have rerouted the fibers on AP table to remove the fiber rolls out of the AP table.  I removed the fibers one by one from both ends - from the 1x16 splitter and from the AP table - keeping the fiber roll intact, and then connected it in reverse way, i.e. the fiber end which was on AP table now is connected to the splitter (since length of the outside the roll is shorter that side) and the fiber end connected to splitter is now rerouted on AP table.

We need to keep the fibers in such a fashion so that no sharp bending occurs anywhere, and also it does not get strained due to its weight, particularly near the 1x16 splitter. Jenne suggested to use a plastic box over the splitter rack to keep the fiber rolls for time-being. We discussed a lot how this can be done nicely; in future we may use array of hooks, Koji suggested to use cable hangers and to tie the rolls using more than one hanging point, Jenne suggested to use the bottom shelf of the rack or to use one plastic box with holes. We tried to make holes on the plastic box using drill, but it developed crack on the box. So ultimately I used the opened box only and put it over the rack.

The corresponding photographs are attached herewith.

Tomorrow we will reroute the fibers for POX table.  

  8505   Mon Apr 29 15:53:49 2013 AlbertUpdateoptical tablesOptical Table Boxes

Quote:

 

 Albert, our new undergrad work force received 40m specific- basic safety training last week. Please read and sign 40m procedures booklet.

 These are the tentative box placements. Roughly. I don't actually have the box finalized yet, but the box should be around that size.

AS1.jpg AS1

BSPRM.jpgBSPRM

ETMX.jpgETMX

ETMY.jpgETMY

MC2.jpgMC2

POX.jpgPOX

POY.jpgPOY

 

How do I perspective ._.

  8504   Mon Apr 29 15:35:31 2013 Max HortonUpdateSummary PagesImporting New Code

 

 I installed the new version of LAL on allegra.  I don't think it has interfered with the existing version, but if anyone has problems, let me know.  The old version on allegra was 6.9.1, but the new code uses 6.10.0.1.  To use it, add . /opt/lscsoft/lal/etc/lal-user-ench.sh to the end of the .bashrc file (this is the simplest way, since it will automatically pull the new version).

I am having a little trouble getting some other unmet dependencies for the summary pages such as the new lalframe, etc.  But I am working on it.

Once I get it working on allegra and know that I can get it without messing up current versions of lal, I will do this again on megatron so I can test and edit the new version of the summary pages.

  8503   Mon Apr 29 08:22:57 2013 SteveUpdateSAFETYsafety training

 

 Albert, our new undergrad work force received 40m specific- basic safety training last week. Please read and sign 40m procedures booklet.

  8502   Mon Apr 29 08:17:40 2013 AlbertUpdateoptical tablesOptical Table Toolboxes Update

Quote:

I decided to go see what the electrical tape looks like on the other tools.

These are the tools I felt were necessary to label with tape: (the others don't seem to be terribly important in terms of not interchanging between boxes)

photo.JPG

On another note I'm not sure why electrical tape can't be used on the Allen Wrenches too.

I also plan on ordering smaller flash lights for each table (this one is bulky and unwieldy), and filling in the gaps of the Allen Wrench sets as soon as I get the go-ahead.

 

  8501   Sat Apr 27 00:29:40 2013 KojiUpdateLSCLSCoffset script fixed

Prior to the locking trials...

scripts/LSC/LSCoffset script had behaved peculiarly:

This script spawns LSC/offset3 in order to remove the dark offset from the channels.
How ever the offsets had been nulled every other PDs
(i.e. The offsets REFL11 I&Q were nulled.
The offsets REFL33 I&Q had been left untouched
The offset REFL55 I&Q had been nulled
and so on.)

I found that the script run many instances of "offset3" scripts in background.
It seemed that tdsavg did not like too many averaging channels at once.

So the "&"s in the LSCoffsets were removed and now the script runs much more slowly,
but works for all of the PDs listed.

I think I have never seen the offsets in REFL33 and REFL165 nulled down to this level before.

  8500   Sat Apr 27 00:21:06 2013 KojiUpdateLSCLocking activity on Apr 26th

When I talked with Den via phone, he recommended to use the trigger and normalization with POP110I.
So I decided to try this approach. Also I investigated how the REFL33 signals are useful.

I could find the state where the PRMI(sb) locks regularly, although the lock is ~1min at most.

PRCL: REFL33I
whitening gain 30dB, -14.0deg (finely tuned in lock)
-> x1.0 -> Triggered by POP110I (20up, 1down)
-> Normalized by POP110I x0.04
-> Gain 0.2~0.12 FM3, 4, 5, 6 always on, no triggered FMs
-> PRM

MICH: REFL33Q
whitening gain 30dB, -14.0deg (finely tuned in lock)
-> x1.0 -> Triggered by POP110I (20up, 1down)
-> Normalized by POP110I x0.04
-> Gain -20 FM4, 5 always on, no triggered FM
-> ITMX (-1.0) and ITMY (+1.0)

I needed to tune the phase very precisely to reach this state. Also the alignment of the michelson and PRM
was very crtiical to acquire the lock.

Later in the same night I was plagued by PRM alignment drift. It seems that the PRM alignment is bistable or
slightly drifting in pitch. I had to align PRM continuously. When the PRMI is locked, the alignment fluctuation
was mainly in yaw. This was as people commented before.

  8499   Fri Apr 26 21:38:06 2013 JenneConfigurationRF SystemPD frequency response

I was sad to see that there wasn't a photo of the POX situation after the fiber work was done on Thursday.

Also, I was out looking at something else, and noticed that the fibers aren't in a very good/safe place from the POX table over to your splitter.  Getting to the POX table is certainly more tricky than the AP table, since the fiber splitter is right next to the AP table, but we should go back and try to make sure the fibers to the more distant tables are laid in a nice, safe way.

Is there a reason that we're not using the clear plastic tubing that Eric bought to put the fibers into?  It seems like that would help a lot in keeping the fibers safe.

I took a few photos of the things that I'm sad about:

1. We should not be keeping fibers on the floor in an area where they can be stepped on.  This will be fixed (I hope) as part of putting the extra coiled length over by the splitter.

IMG_0498.JPG

2. Again, in an area where we semi-regularly walk, the fibers should not be a tripping hazard.  Behind the table legs (rather than under the middle of the table) is safer, and will help tuck them out of the way.

IMG_0499.JPG

3.  It's not obvious when we're pumped down, but we remove the access connector (top right side of this photo), and need to walk in this area.  I can pretty much guarantee that within 1 day of the next time we vent, these fibers will be stepped on, tripped over, and broken if they are not moved to a different location.  I'm not yet sure what the best way to route these fibers is, but this is not it. 

IMG_0500.JPG

Riju, since Eric will be away next week, please let one of us "40m Regulars" know when you plan to come over (at least a few hours ahead of time), and we can give you a hand in protecting these fibers a little bit better.  Thanks!

  8498   Fri Apr 26 20:43:51 2013 JenneUpdateLSCRemeasuring the Schnupp asymmetry

[Jenne, Annalisa, with guidance from Koji]

We took data to remeasure the Schnupp asymmetry, using the Valera method that Jamie described in elog 4821

1  First, we locked the arms each with their PO(X,Y) signals, to get the alignment of each arm. 

2.  Then, we locked the Xarm with AS55I (Yarm optics, and PRM very misaligned, more than the misalign script).  Since AS55 was saturating, I changed the analog gain from 24dB to 21dB. (After work was completed, the analog gain was put back to the nominal 24dB for both I&Q.)

3.  We set up the Lockin similar to Jamie's description, with a few differences.  We used the same f = 103.1313, but used ampl=10cts.  Sin and cos gain were each 100.  We changed the lowpass filter from 0.1Hz to 0.05Hz (so each measurement had a settling time of at least 20sec).  We were using LSC-Lockin4, so the Lockin matrix was set so Lockin4 was reading from AS55Q, and the LSC output matrix was such that we were actuating on the ETM (X, then Y when we switched arms later).

4.  By hand, we roughly found the zero crossing of the lockin-q output (which corresponded also to zero of the lockin-I, since this is the place where all of the PDH signal was in AS55I, and the lockin was reading AS55Q). 

5.  We took points separated by 0.2 degrees, plus and minus 1 degree from the zero-crossing phase we had found (i.e., for the Xarm, we roughly found the zero crossing at -14.39 deg, so took data from -15.39 to -13.39degrees).  For each phase, we took 5 measurements (using ezcaread), at least 20 seconds apart.  After moving the phase, we waited at least ~40 seconds (watching the lockin outputs on striptool, they had completely settled after 30 or 40 seconds).

6.  We then repeated steps 2, 4 and 5 for the Y arm.  The lockin setup didn't change, except that now we actuate on ETMY.

We did a quick estimate calculation, from our rough zero-crossings to get a rough measurement of the Schnupp asymmetry.  DeltaPhi = (-14.39 -   -19.79) = 5.40 . This gives us (using F_sideband = 5*11066134, the current 11MHz marconi freq) a rough Schnupp asymmetry of 4 cm. 

Analysis to follow.

EDIT, JCD:  The Xarm gain at this time was -0.160, and the Yarm gain was -0.170

  8497   Fri Apr 26 17:08:42 2013 RijuConfiguration PD frequency response

Quote:

No.... what I told was to put the roll next to the splitter, not on the table.
The table area is more precious than the rack space.

Koji> The slack of the fibers should be nicely rolled and put together at the splitter side.

 Ok, will do it on the coming week.

  8496   Fri Apr 26 15:50:48 2013 Max HortonUpdateSummary PagesImporting New Code

 

I am following the instructions here:

https://www.lsc-group.phys.uwm.edu/daswg/docs/howto/lal-install.html#test

But there as an error when I run the ./00boot command near the beginning.  I have asked Duncan Macleod about this and am waiting to hear back.

For now, I am putting things into /home/controls on allegra.  My understanding is that this is not shared, so I don't have a chance of messing up anyone else's work.  I have been moving slow and being extra cautious about what I do because I don't want to accidentally nuke anything.

  8495   Fri Apr 26 10:50:07 2013 AnnalisaUpdateABSLATF laser on PSL

The ATF NPRO auxiliary laser has been moved on the PSL table. All the optics for beat note measurement are in place and alignment has been done.

The setup for this measurement is the same as described in elog 8333.

  8494   Thu Apr 25 20:48:48 2013 KojiUpdateASSDen fixed the Yarm ASS scripts

I contacted Den about malfunctioning of the Yarm ASS.

He found the scripts were modified during the attempt to make it available for Xarm (cf. a related elog entry)
So far, he could manage to make the current scripts being modified to run.
A striptool file is still missing but this is what we can handle locally.

I thank Den for the remote caring of the issue despite the limited network bandwidth.

  8493   Thu Apr 25 18:58:06 2013 KojiConfiguration PD frequency response

No.... what I told was to put the roll next to the splitter, not on the table.
The table area is more precious than the rack space.

Koji> The slack of the fibers should be nicely rolled and put together at the splitter side.

  8492   Thu Apr 25 17:56:28 2013 RijuConfiguration PD frequency response

 [Eric, Riju]

Today we have routed the fibers from 1x16 fiber splitter to POX table for POX11 PD and POP55 PD. Also we labeled the fibers on AP table, they have been fixed on the table. The photo of the table after work is attached here. We will do it for POX table tomorrow. 

  8491   Thu Apr 25 10:19:10 2013 KojiSummaryLSCLocking activity on Apr 24th

Last night I worked on the several locking configurations:

General preparations / AS table inspection

- The AS beam looked clipped. I went to the AP table and confirmed this is a clipping in the chamber.
  This may be fixed by the invacuum PZTs.

Modulation frequency tuning

RFPD Mon of the MC demodulator was check with the RF analyzer. Minimized the 25.8MHz (=55.3-29.5MHz) peak by changing the marconi freq.
This changed the modulation freq from 11.066147MHz to 11.066134MHz. This corresponds to the change of the MC round-trip length from
27.090952m to 27.090984m (32um longer).

Michelson tests

- I wonder why I could not see good Michelson signal at REFL ports.

- I roughly aligned the Michelson. On the AP table, the RF analyzer was connected to the REFL11 RF output.
  By using "MAX HOLD" function of the analyzer, I determined that the maximum output of the 11.07MHz peak
  was -61.5dBm.

- I went to the demodboard rack. I injected -61dBm from DS345 into the RFEL11 demodboard. This produced
  clean sinusoidal wave with the amplitude of 4 count. The whitening gain was 0dB.

- The output from the PD cable was -64.0dBm. So there is ~2.5dB loss in the cable. Despite this noise, the demodulation
  system should be sufficiently low noise. i.e. the issue is optical

- The Michelson was locked with AS55Q. And the REFL11 error signals were checked.Fringe like feature was there.
  This suggested the scattering from the misaligned PRM. The PRM was further misaligned. Then some reasonable
  (yet still noisy) Michelson signal appeared. (Usual misaligned PRM is not at the right place)

  Q. How much scattering noise (spurious cavity between PRM and the input optics) do we have when the PRM is aligned?
  Q. Where should we put the glass beam dumps in the input optics?
  Q. Can we prepare "safe" misaligned place for the PRM with the beam dump?

- The Michelson was locked with REFL11Q. From the transfer function measurement, the gain difference between AS55Q (whitening gain 24dB)
  and REFL11Q was 32dB. The whitening gain was 0dB. In fact I could not lock the Michelson with the whitening gain 33dB (saturation???)
  The element in the Input matrix was 1, The gain of the servo was +100. BS was actuated.

Coupled cavity tests

- At least REFL11 is producing reasonable signals. So what about the other REFL ports? The Michelson signals in the other frequencies
  were invisible. So I decided to use three-mirror coupled cavity with the loss PRC.

- Aligned X arm, Misaligned ETMX, ITMY. Aligned PRM.

- Locked the PRM-ITMX cavity with REFL11 and REFL33.

- Aligned ETMX. If I use REFL11I for the PRC locking, I could not lock the coupled cavity. But I could with REFL33I.
  This is somewhat familiar to me as this is the usual feature of the 3f signal.

- The coupled cavity could be locked "forever". To realize this I needed to tweak the normalization factor from 1.0 to 1.6.
  Q. How does the coupled cavity change the response of the cavity? Can we compensate it by something?
  Q. Measure open loop transfer functions to check if there is any issue in the servo shapes.

- Transmission during the lock is 3.2 while the nominal TRX with PRM misaligned was 0.93.
  This corresponds to power recycling gain of 0.17.

 - X arm:

    - Source: POX11I, phase 79.5 deg, whitening gain 36dB
    - Input matrix: POX11I->1.0->XARM, Normalization TRX*1.60
    - XARM servo gain +0.8, actuation ETMX
    - XARM trigger 0.25 up, 0.05 down. XARM Filter trigger untouched.

- PRC: (sideband locking)
    - Source: REFL33I, phase -34.05 deg, whitening gain 30dB
    - Input matrix: REFL33I->1.0->PRCL, Normalization None
    - PRCL servo gain +4.0, actuation PRM
    - PRCL trigger None

- Same test for the Y arm. At the moment ETMY did not have the OPLEV.
  Same level of transmission (~3.3)

 - Y arm:

    - Source: POY11I, phase -61.00 deg, whitening gain 36dB
    - Input matrix: POY11I->1.0->YARM, Normalization TRX*2.1
    - YARM servo gain +0.25, actuation ETMX
    - YARM trigger 0.25 up, 0.05 down. YARM Filter trigger untouched.

- PRC: (sideband locking)
    - same as above

Sideband PRMI attempt

    - Now I got some kind of confidence on the REFL33 signal.
    - So I tried to get any stable setup for sb PRMI, then to find any reasonable MICH signals anywhere else than AS55Q.
    - With REFL33I(PRCL) & AS55Q(MICH), I got maximum ~10sec lock. It regularly locked. It was enough long to check
      the spectrum on DTT. But it was not enough long to find anything about the MICH signals at the REFL ports.

    - I tried REFL33Q for MICH. The lock was even shorter but could lock for 1~2 sec.

    Q. What is the cause of the lock loss? I did not see too much angluar fluctuation. The actuation was also quiet (below 10000).

- PRCL: (sideband locking)
    - Same as above except for
      - the PRCL servo gain +0.05, No limitter at the servo output.
      - Trigger POP22I (low pass filtered by LP10) 20 up, 3 down

- MICH:
    - AS55Q -24.125 24dB -> x1.0 -> MICH -0.7, No limitter -> ITMX/Y differential
    or
    - REFL33Q -34.05dB -> x2.0 -> MICH same as above
    - For both case, trigger POP22I (low pass filtered by LP10) 20 up, 3 down

 

At this point Jenne came back from dinner. Explained what I did and handed over the IFO.

  8490   Thu Apr 25 04:10:09 2013 JenneUpdateLockingMICH_CTRL drifting away??

Koji is elogging separately of his exploration of different configurations.  The lock stretch that I'm looking at here uses the same parameters as Koji had for PRMI sb lock, using AS55Q for MICH and REFL33I for PRCL, with MICH gain of -0.8 and PRCL gain of 0.05 .

All of these plots are the same few second lock stretch, with different zooming.  Jamie's super-sweet getdata python script only accepts integers for the start time and duration parameters, so lots of this zooming happened by hand, but I tried to always keep the time axis aligned within each screenshot.  Sometimes the plot axis labels say differently, but they're lying to you.

Plot 1:  gps start time is 1050915916, duration = 6 seconds.  Overall view of the lock stretch.

1050915916-6.png

Plot 2:  gps start time is 1050915921, duration = 1 second.  We're looking at the lockloss that happens at the left side of the plots.

1050915921-1.png

Plot 3:  zoomed in (along the time-axis) version of plot 2, so much shorter time duration.  Some zooming on y-axes.

1050915921-zoom.png

Plot 4:  zoomed in (along y-axes) version of plot 2.

1050915921-1-zoom.png

It seems to me from these plots that maybe MICH CTRL is drifting away?  It seems like we lose the MICH lock, and that destroys the whole thing. 

Koji made some comments to me earlier, regarding his work this evening, that the MICH signal quality is poor in general, and that we should calculate/think about changing our schnupp asymmetry. 

  8489   Thu Apr 25 03:35:28 2013 JenneUpdateLockingAngular motion does not explain RIN

Den made a nice elog about the PRMI RIN that we see a few weeks ago:  8464.  The RIN that we're seeing is typically about ~30%.  The question at hand is: what is causing this power fluctuation, and more specifically, is it the angular motion of the mirrors?

I find that no, the angular motion that we see does not explain the RIN that we see.

In the attached Mathematica notebook, I calculate the power lost due to angular misalignments of one or more mirrors.  (Math comes from Appendix A of Keita's thesis.)

From calibrated oplev spectra, our mirrors are moving about 1 microradian (RMS, which is dominated by low frequencies).  From a super sophisticated "draw on the TV, then measure" method (details below), I have estimated that the maximum static misalignment that we're seeing is about 2 microradians.

With all of this, I find that for a g-parameter of 0.94, the power lost due to misalignments should, at maximum, be 0.6%.  I need a g-parameter of 0.995 to get a power loss of 23%.  Alternatively, if I take the derivative of the power coupling function, to find the static misalignment at the steepest slope of the curve (and thus, the place where any AC misalignment would have the most effect), for 1urad of AC misalignment, I get 40% power loss. 

So, in order for the AC angular motion that we see to explain the RIN that we see, either our mirrors are very, very misaligned (so much so that we couldn't really be locking), or our cavity is much closer to unstable than expected from Jamie's calculations.  Since both of these cases (static misalignment or incorrect g-parameter calculation) have to be taken to extremes before they approximate the RIN that we see, I do not think that this power loss is due to angular fluctuations.

This means that we have to think of another potential cause for this RIN that we're seeing.

Details on the "draw on TV and measure" technique for determining static cavity misalignments:  Looking at the POP camera view, with the PRM significantly misaligned, I traced the straight-through beam spot.  I then restored the PRM, and during several momentary locks, I traced the beam spot, which I took to be the saturated area of the camera.  The idea here is that the straight-through beam represents the incident beam axis, while the locked beam represents the cavity axis.  I'm assuming that the camera image plane is at the face of PR2. I approximately found the center of each of my tracings, and found them to be ~1/4 inch apart.  I also measured the "spot size" of the sideband-locked PRMI, and found it to be ~3.5 inches.  So, very roughly, the ratio of (distance between spots)/(size of beam) is ~0.07. This corresponds to a static misalignment of either the ITM or the PRM of ~2urad, rounding up. (I use the Jamie's calculated g-parameters from elog 8316, the case of flipped PR2, tangential = 0.94 to calculate the effective RoC of the PRM). 

  8488   Thu Apr 25 00:59:37 2013 RijuUpdateRF SystemPD frequency response

Quote:

I think you have the splitter that splits the RF signal from the network analyzer in the wrong place. 

Usually you split the signal immediately after the RF Out, so that half of the signal goes to the A-input of the Analyzer, and the other half goes to your controller (here, the laser diode controller).  Then you would take the output of your controller and go straight to the actual laser diode, with no splitting in this path.

 Here our device under test is the photodiode. So for the reference I wanted to retain the response of the laser diode controller. Otherwise I have to consider the transfer function of that LDC too. I may check both the options at the time of experiment.

Thanks

  8487   Wed Apr 24 18:51:12 2013 KojiConfigurationoptical tablesPD frequency response

The fibers should be routed beneath the electrical cables.
They should be fixed on the table for strain relieving.
The slack of the fibers should be nicely rolled and put together at the splitter side.

These are expected to be done next time when the fiber team work around the table.

We also expect to have the table photo every time the work of the day is finished.

  8486   Wed Apr 24 15:27:59 2013 ManasaUpdate40m UpgradingEndtable upgrade for auxiliary green laser : progress: New layout

Layout that will be improved upon over the next few days.

Things that need to be updated:
1. Waist size at all optics
2. Beam size at detectors and choice of lenses
3. IPANG & green PD proposed positions

 ETMY_endtable_New.png

  8485   Wed Apr 24 14:36:06 2013 JenneUpdateRF SystemPD frequency response

I think you have the splitter that splits the RF signal from the network analyzer in the wrong place. 

Usually you split the signal immediately after the RF Out, so that half of the signal goes to the A-input of the Analyzer, and the other half goes to your controller (here, the laser diode controller).  Then you would take the output of your controller and go straight to the actual laser diode, with no splitting in this path.

  8484   Wed Apr 24 14:24:40 2013 RijuUpdate PD frequency response

 Here I am attaching the first schematic diagram of the PD frequency response set-up, I will keep updating it with relevant informations with the progress of the work.

Description: Our objective is to set-up one simultaneous transfer-function measurement system for all the RF-PDs present in 40m lab. A diode laser will be used to illuminate the PDs. The diode laser output will be divided by 1x16 fiber splitter and will be sent to all the PDs through single-mode fiber. The transfer function of the PDs will be measured using network analyzer(Agilent 4395A). The output of the PDs will be fed to network analyzer via one RF-switch. The diode laser will be controlled by the controller ILX LDC 3744C. The scanning frequency signal will be fed to this controller from network analyzer through its external modulation port. The output of the controller will be splitted  into two parts: one will go to laser diode and the other will be used as reference signal for network analyzer.

 

 

  8483   Wed Apr 24 14:20:49 2013 KojiUpdateCDSFE Web view not updated?

The FE web view seems not up-to-date, does it? ( maybe for a year)

https://nodus.ligo.caltech.edu:30889/FE/c1mcs_slwebview_files/index.html

  8482   Wed Apr 24 00:44:33 2013 KojiUpdateSUSPRMI locked, ITMX pitch OpLev ringing up

I tried to reproduce the locking situation described in this entry tonight.
The momentary lock was regularly seen but there was no stable lock.

I wonder why the actuators are always saturated. The feedback signals have the dominant component at ~400Hz.

It would also be nice if the servos have some immunity to gain fluctuation.

I didn't check how the situation of the AP table is. I'll look into some details tomorrow.

  8481   Wed Apr 24 00:42:07 2013 JenneUpdateSUSPRMI locked, ITMX pitch OpLev ringing up

Koji is working on PRMI locking, and while he was doing that I glanced at the oplevs' spectra for the ITMs and PRM.

I found that when the PRMI was locked (for only 1 second or so max lock time) on the 55MHz sideband, and the error signals show a big peak around 400Hz (definitely audible in the control room), the only OpLev that I see a similar peak in is ITMX pitch. 

In the plot below, I have grabbed a time when the PRMI was flashing as the black reference traces, and then a time when the PRMI was locked as the active traces.  You can see that there is a similar peak in both REFL55I and ITMX_OL_PIT when the cavity is locked. 

PRMI_locked_ITMXpitOpLevRingingUp.pdf

  8480   Tue Apr 23 22:59:05 2013 ranaConfigurationSUSOptical Lever Gains normalized

Due to the recent addition of cal factors in the OL error points, the OLPIT_GAIN and OLYAW_GAIN have been reduce to tiny numbers (e.g. 0.002).

Since our MEDM only shows 3 digits past the decimal point by default, it makes more sense to have the gains around 1.

So I reduced the gains in all of the FM1 filters from 1000 to 1 and multiplied the GAIN values by 1000 (using ezcastep) to compensate.

All of the active optics seem to be behaving as before. Haven't tested ETMs or SRM yet.

  8479   Tue Apr 23 22:10:54 2013 ranaUpdateComputersNancy

controls@rosalba:/users/rana/docs 0$ svn resolve --accept working nancy
Resolved conflicted state of 'nancy'

  8478   Tue Apr 23 16:31:13 2013 EricConfiguration PD frequency response

[Eric, Riju]

Summary: Routing Fibers on AP table for Photo Diode Frequency Response Measurement System

Objective: We are to set-up one simultaneous transfer-function measurement system for all the RF-PDs present in 40m lab. A diode laser output is to be divided by 1x16 fiber splitter and to be sent to all the PDs through single-mode fiber. The transfer function of the PDs will be measured using network analyzer. The output of the PDs will be fed to network analyzer via one RF-switch.

Work Done So Far: We routed the fibers on AP table. Fibers from RF PDS - namely  MC REFL PD, AS55, REFL11, REFL33, REFL55, REFL165, have been connected to the 1x16 fiber splitter. All the cables are lying on the table now, so they are not blocking any beam.

We will soon upload the schematic diagram of the set up.

 

Missing Component: Digital Fiber Power Meter, Thorlab PM20C

 

 

  8477   Tue Apr 23 16:17:45 2013 DenUpdatePEMseismometer isolation kit in place

Quote:

 

 The carpenter shop finished the installation of the 3 granite bases.Rapid Set Cement All high strength non-shrink grout was used.

 Compressive strength  3000 PSI at 1 hour and 9000 PSI at day 28 The janitor is still cleaning up after them at the south end.

The  soft silicon gas kits are working well with the SS can.  Den is making  the adaptor plate drawing for the feedthrough.

 To put everything in one place I add a final drawing of the base to this elog.

 Next time we continue with wiring and putting temperature and pressure sensors inside the box. Connector support plate drawing is attached. We'll have sensors inside the kit with STS-2 or Trillium as their connector is small enough (19 pin vs 26 pin for Guralps) that we can put an additional 4 pin lemo connecor (2 pins for each sensor). I think EGG.0B.304.CLL is good for this application. Temperature and pressure sensor we can by from omega.

  8476   Tue Apr 23 15:02:19 2013 Max HortonUpdateSummary PagesImporting New Code

Duncan Macleod (original author of summary pages) has an updated version that I would like to import and work on.  The code and installation instructions are found below.

I am not sure where we want to host this.  I could put it in a new folder in /users/public_html/  on megatron, for example.  Duncan appears to have just included the summary page code in the pylal repository.  Should I reimport the whole repository?  I'm not sure if this will mess up other things on megatron that use pylal.  I am working on talking to Rana and Jamie to see what is best.

http://www.lsc-group.phys.uwm.edu/cgit/lalsuite/tree/pylal/bin/pylal_summary_page
https://www.lsc-group.phys.uwm.edu/daswg/docs/howto/lal-install.html
  8475   Tue Apr 23 15:00:20 2013 JenneUpdate40m Upgrading4 pins enough?

Quote:

Are 4 of these spring loaded pins enough?  I'm not sure how one pin can hold 2 lids at each point.  It seems like we need 8 pins.

 Steve has explained to me that the pins will go in between the 2 lids, with a big washer, so that one pin holds both lids at the same time.  4 is the right number.

  8474   Mon Apr 22 20:17:05 2013 CharlesUpdateISSNew Servo w/switching filters

 

In my previous post here, a new servo design was discussed. Although the exact design used will depend on the particular noise requirements for the 40m and the Bridge Labs (requirements will be considered separately for each application), I still have to yet to see those formalized. Despite this, I have been simulating an example servo circuit with three switchable stages. The design can be found at: New Servo.

Essentially, this circuit consists of three unity gain buffers that can be switched into different filtering states. Attached is a plot of the transfer function of this particular circuit with successive stages turned on. The curve (0) corresponds to all of the filters being switched off, so the total behavior is that of a unity gain buffer. The curve (1) corresponds to the first stage being turned on with the 2nd and 3rd still acting as unity gain buffers. This first state has a gain of ~80 dB at DC and a pole at ~10 Hz which sets the unity gain crossing at ~100 kHz. The curves (2) and (3) correspond to the second and third stage being turned on, respectively. Each of these stages has a pole at DC (i.e. ~infinite gain) and a zero at 10^4 Hz. For f > 10^4 Hz, these stages have gain ~ 1, as we can see in the transfer function below.

I have also performed some noise analysis of this circuit. Attached are a few plots produced by LISO showing the resistor and op-amp noise separately (it was too cluttered on one plot) at the output node of the servo. Both of these plots have a "Sum Noise" trace, which is the sum for every circuit element and is thus identical between plots. The third noise spectrum included is simply the noise at the output referenced to the input with the previously computed transfer function. I'm not sure if there is a simple method embedded in LISO to reference the noise at the output node to the input, but it should be as simple as numerically dividing the noise spectrum by the transfer function between input and output. 

Next, I will be attempting time-dependent simulations of this simple circuit using delayed switches instead of manually controlled ones.

  8473   Mon Apr 22 19:48:56 2013 JenneUpdate40m Upgrading4 pins enough?

Are 4 of these spring loaded pins enough?  I'm not sure how one pin can hold 2 lids at each point.  It seems like we need 8 pins.

  8472   Mon Apr 22 17:43:09 2013 SteveUpdate40m Upgrading ETMY optical table & enclosure is ready for optics

Quote:

Quote:

 Enclosure is at the east end. It has it's bottom o-ring in place. It will be ready for optics tomorrow around 5pm

I have to shim out the enclosure, finish leveling the table and cut surgical tubing O-ring for the top.

 

 Glued surgical latex tubing with super glue into O-ring shape. The existing in place tubing K-100, OD 0.125" (actual size 0.140"), wall 0.031", ID 0.062".

I have just found out that tolerances on tubing OD are + - 0.026" by the manufacturer. I'm getting larger tubing for better fit.

The table is ready for optics.

Things left to do:

1, finalize o-ring size  2, finish cable feedthrough  3, finalize window connection 4, IR-Thermashield strips for bridge sides 5, replace bridge support post with solid one

 

 I'm working on to improve the quality of the enclosure.

The short comings are: more cable feedthroughs needed, latches to anchor top covers air tight and posts to support bending bridges.

Red triangles are compression latches at 10 places to hold the top air tight on surgical tubing

Green lines represent 4  posts of Al 1" OD to support the covers and maximize their eigenfrequencies.

Black crosses are 4 spring loaded push-bottom quick release pins to anchor the top covers to the bridges. This connection will  not be air tight.

(quarter-turn wing head fastener have the same problem) I'm thinking of some solution to minimize the leak.    

Violet _ steel plate (1" wide, 15" long, 0.125" thick)  between the two posts will anchor the quick release pins and make bridge rigid.

Blue rectangle is an other cable feedthrough exiting on the chamber side.

Planning to substitute window with soft - air tight ( Aluminized thin wall hose )  connection to vacuum view port where white circle is representing the Al adaptor ring.

Updated after Wednesday meeting 4-24-2013

 

  8471   Mon Apr 22 17:06:42 2013 ranaSummaryIOOMC locked/aligned. MC WFS offloading by ezcaservo

 

 Why use the PSL beam as a reference? Don't we want to keep the MC pointing in a good direction through the Faraday instead???

  8470   Mon Apr 22 12:03:58 2013 KojiUpdatePSLPMC aligned too

PMC aligned. C1:PSL-PMC-PMCTRANSPD improved from 0.72ish to 0.835ish.

  8469   Mon Apr 22 11:46:09 2013 KojiSummaryIOOMC locked/aligned. MC WFS offloading by ezcaservo

Еру ьс шы тщц дщслув фтв фдшптувю

Фдыщ ш кфт еру ащддщцштп ыскшзе ещ щаадщфв еру ЬС ЦАЫ ыукмщю

I blame Den for russian keyboard installation on the control machines.

ezcaservo -r 'C1:SUS-MC2_ASCPIT_OUT16' -g '0.00001' -t 60 C1:SUS-MC2_PIT_COMM&
ezcaservo -r 'C1:SUS-MC2_ASCYAW_OUT16' -g '0.00001' -t 60 C1:SUS-MC2_YAW_COMM&
ezcaservo -r 'C1:SUS-MC1_ASCPIT_OUT16' -g '0.00001' -t 60 C1:SUS-MC1_PIT_COMM&
ezcaservo -r 'C1:SUS-MC1_ASCYAW_OUT16' -g '0.00001' -t 60 C1:SUS-MC1_YAW_COMM&
ezcaservo -r 'C1:SUS-MC3_ASCPIT_OUT16' -g '0.00001' -t 60 C1:SUS-MC3_PIT_COMM&
ezcaservo -r 'C1:SUS-MC3_ASCYAW_OUT16' -g '0.00001' -t 60 C1:SUS-MC3_YAW_COMM&

ELOG V3.1.3-